OpenStudy (anonymous):

Choose the correct graph of the equation. y - 2 = 2/3(x - 4)

OpenStudy (anonymous):

Here are the options

OpenStudy (anonymous):

First, multiply out 2/3 so you have y-2=2/3x -8/3 then add 2 to both sides y=2/3x-2/3 look for a graph that goes up two units for every three units to the right it goes and intercepts the y-axis at -2/3

OpenStudy (anonymous):

OK Thanks! That would be the first option :)

OpenStudy (anonymous):

Actually, it would be the third option @JoStudy if you look at where the line meets the y-axis it is between -1/2 and -1 so that would be -2/3

OpenStudy (anonymous):

OH OK! Now I get it. I wasn't reading -2/3 as a fraction I was reading it as -2! Thanks!